OpenStudy (anonymous):

What is the simplified form of (y^2-6y+8)/(y^2-7y+12)? i need some HELP!!!

OpenStudy (cwrw238):

can you factor the two quadratics on top and bottom?

OpenStudy (cwrw238):

they both factor to form ( y +/1 a) (y +/- b)

OpenStudy (anonymous):

yes thats what i got

OpenStudy (cwrw238):

y^2-6y+8) = (y - 4)(y - 2) and y^2-7y+12 = (y - 4)(y - 3) so (y - 4)(y - 2) --------- (y - 4)(y - 3) simplifies to what?

OpenStudy (anonymous):

(y-2)/(y-3)

OpenStudy (cwrw238):

what factor is common to both top and bottom of the fraction?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

y-4 / y-4 is of course = 1

OpenStudy (anonymous):

thankyou soo much for clearing that up for me @cwrw238 @MathLegend

OpenStudy (cwrw238):

no probs - glad to be of help

OpenStudy (cwrw238):

(the bills in the post - lol )

OpenStudy (anonymous):

lol thanks a millon